# Importance Sampling-Guided Meta-Training for Intelligent Agents in Highly Interactive Environments

### v1

@@ IV-A Synthetic Naturalistic Distribution Example
In this example, we let $p_{\text{naturalistic}}=N(1.5,0.5^{2})$ and $p_{\text{training}}=N(0.5,0.5^{2})$, a setting that allows the ego policy to be trained on a larger proportion of dangerous scenarios (lower $\beta$ values) compared to what would have been observed from the naturalistic distribution.
We train four ego policies: GEP and GIS (using $p_{\text{uniform}}$), NEP (using $p_{\text{naturalistic}}$), and CEIS (using $p_{\text{training}}$) with $K=1$ (no iterative training). We use $p_{\text{naturalistic}}$ to generate $\beta$, simulating naturalistically random social vehicle behavior. Each policy is evaluated over a total of $N_{s}=$5,000$$ samples. This experiment is meant to study whether a different training distribution leads to different performance results under naturalistic evaluation conditions. The GEP is used as a baseline as in the original paper [lee2023robust].
In Table I, we report the number of successes (the ego vehicle passed the intersection without collision), collisions, and time-out (the ego vehicle did not cross the intersection to avoid collisions). We also tested 10 other combinations of $\mu\in\{-0.5,0.5,1.5,2.5\}$ and $\sigma\in\{0.5,0.75,1.0\}$ for Gaussian $p_{\text{training}}=N(\mu,\sigma^{2})$ (not running CE and mixture model in Algorithm 1) and summarized the result in Table II.
TABLE I: Ego policies evaluation for the synthetic example
Ego Policy Training Distribution Success $\uparrow$ Collision $\downarrow$ Time Out $\downarrow$
GEP $U(-1,3)$ 0.660 $\pm$ 0.473 0.334 $\pm$ 0.472 0.006 $\pm$ 0.077
GIS $U(-1,3)$ 0.881 $\pm$ 0.337 0.107 $\pm$ 0.311 0.103 $\pm$ 0.131
NEP $N(1.5,0.5^{2})$ 0.681 $\pm$ 0.466 0.119 $\pm$ 0.137 0.200 $\pm$ 0.320
CEIS $N(0.5,0.5^{2})$ 0.811 $\pm$ 0.392 0.108 $\pm$ 0.393 0.081 $\pm$ 0.311
@@
Std Mean Metric $\mu=-0.5$ $\mu=0.5$ $\mu=1.5$ $\mu=2.5$
$\sigma=0.5$ Success $\uparrow$ 0.761 $\pm$ 0.424 0.811 $\pm$ 0.392 0.681 $\pm$ 0.466 0.701 $\pm$ 0.400
Collision $\downarrow$ 0.037 $\pm$ 0.189 0.108 $\pm$ 0.393 0.119 $\pm$ 0.137 0.218 $\pm$ 0.413
Time Out $\downarrow$ 0.202 $\pm$ 0.110 0.081 $\pm$ 0.311 0.200 $\pm$ 0.320 0.031 $\pm$ 0.081
$\sigma=0.75$ Success $\uparrow$ 0.774 $\pm$ 0.129 0.786 $\pm$ 0.277 0.765 $\pm$ 0.312 0.783 $\pm$ 0.331
Collision $\downarrow$ 0.124 $\pm$ 0.153 0.168 $\pm$ 0.252 0.151 $\pm$ 0.459 0.196 $\pm$ 0.308
Time Out $\downarrow$ 0.102 $\pm$ 0.141 0.046 $\pm$ 0.125 0.084 $\pm$ 0.182 0.020 $\pm$ 0.313
$\sigma=1.0$ Success $\uparrow$ 0.679 $\pm$ 0.485 0.703 $\pm$ 0.398 0.681 $\pm$ 0.466 0.689 $\pm$ 0.314
Collision $\downarrow$ 0.218 $\pm$ 0.375 0.190 $\pm$ 0.392 0.214 $\pm$ 0.464 0.193 $\pm$ 0.291
Time Out $\downarrow$ 0.103 $\pm$ 0.298 0.107 $\pm$ 0.284 0.105 $\pm$ 0.271 0.118 $\pm$ 0.233
@@ IV-B Real-World Naturalistic Distribution
[…]
@@
With this naturalistic distribution, we then compare the four benchmarks (GEP, GIS, NEP, and CEIS). In this case, GEP and GIS use $U(-1,3)$, similar to the previous example. NEP uses the fitted KDE at T-Intersection #1 as $p_{\text{training}}$. CEIS uses the best-fitted Gaussian from the fitted KDE at T-Intersection #2 (which in this case turns out to be $N(1.8,0.192^{2})$ as $p_{0}$ in Algorithm 1. Both GEP and NEP are trained only once (since no iteration is needed) while CEIS is trained with $K=3$ (running CE offline to optimize $p_{\text{evaluation}}$ and using GMM approach for the subsequent $p_{k}$’s as shown in Algorithm 1). The result is summarized in Table III. We further project the CEIS performance over more iterations and show the summary in Figure 4.
TABLE III: Ego policies evaluation for the naturalistic example
Ego Policy Training Distribution Success $\uparrow$ Collision $\downarrow$ Time Out $\downarrow$
GEP $U(-1,3)$ $0.757\pm 0.440$ $0.235\pm 0.426$ $\mathbf{0.008\pm 0.106}$
GIS $U(-1,3)$ $0.851\pm 0.380$ $0.128\pm 0.339$ $0.021\pm 0.160$
NEP KDE $0.816\pm 0.395$ $0.081\pm 0.286$ $0.103\pm 0.334$
CEIS, $k=1$ $N(1.8,0.5^{2})$ $0.834\pm 0.382$ $0.108\pm 0.318$ $0.058\pm 0.257$
CEIS, $k=2$ GMM $0.893\pm 0.320$ $0.075\pm 0.293$ $0.032\pm 0.200$
CEIS, $k=3$ GMM $\mathbf{0.934\pm 0.277}$ $\mathbf{0.033\pm 0.189}$ $0.033\pm 0.207$
@@ V Discussion
We discuss our findings in terms of the CEIS performance over the baselines, its generalization capability especially under naturalistic settings, and its improvement over training iterations.
@@ V-A Dominant IS-based Performance Over the Baselines
In both experiments, the IS-based algorithms (the GEP and NEP) consistently outperformed their counterparts across success rates and collision rates (see Table I and Table III). Their performance is slightly lower in terms of timeout rates. Notably, GIS achieves an 88.0% success rate and 10% collision rate in the synthetic examples, while CEIS achieved an 81.1% success rate and 10.8% in collisions (over +10% improvement in success rates vs non-IS-based approaches). Meanwhile, NEP achieves only 68% and GEP only 66% success rates. The collision rate for GEP and NEP is 33% and 12%, respectively.
These results show the efficiency of the IS-based approach in training more capable and safer autonomous agents (lower collision rates). The only slight downside is in the time-out rate, where the GEP dominates with a 0.6% timeout rate, highlighting its focus on finishing the task (either successfully or not). We argue, however, that in many robotics settings, it is more desirable to have an agent that is not only focusing on just finishing the task (while endangering its platform and surroundings) but rather to have a safe agent that avoids risking catastrophic costs. To that end, the IS-based approach seems to be capable of reasoning about the risk better by considering the likelihood of any risky maneuver when making decisions (see a visual example in Figure 3). This is the strength of IS that we leverage in this work. By evaluating performance across a mix of common and rare cases, IS enables the training to adapt more appropriately to the dynamics of real-world driving scenarios. This methodological advantage translates into an agent that is not only highly effective in common scenarios but also capable of handling unexpected or extreme situations with reasonable performance.
((a)) Initial T-intersection state. Blue shows ego vehicle. Red shows social vehicles.
((b)) Failure example, where merging is initiated too quickly causing a crash.
((c)) Success example, showing the rear vehicle is given sufficient time to react.
Figure 3: Ego vehicle policy in the simulated T-intersection.
@@ V-B Better Results in Naturalistic Settings
This observation is also supported by Table II, where a more concentrated training distribution tends to give higher rewards. Combined with its iterative training scheme, CEIS can improve over time based on the initial successes, suggesting that continuous learning and adaptation are feasible and effective, providing a pathway to even more sophisticated and refined driving policies. This characteristic is crucial for the deployment of autonomous agents in highly interactive driving environments.
((a)) Failing to cross overcrowded intersection.
((b)) Could not mitigate adversarial behavior (e.g. sudden acceleration). Red arrows show the displacements within one timestep.
@@
V-C Trend of Improving Performance Over Time
A longitudinal evaluation of CEIS apparent from Figure 4 reveals a positive trend of incremental performance gains over successive training iterations. Specifically, the success rate improved from 83% in the first iteration to 98% (projection), with a noted decrease in collision rates from 10% to 0.5% in 10 training iterations. This trend illustrates the effectiveness of combining CE methods with a GMM to bridge guided $p_{\text{evaluation}}$ to $p_{\text{training}}$, ensuring balanced and comprehensive training. However, the tapering off of success rates beyond the fourth iteration—from 95% to a seemingly asymptotic 98%—highlights a potential ceiling in performance improvements, suggesting diminishing returns as the training process approaches optimization saturation. This limitation is an essential area for further investigation, suggesting that CEIS might try to address the full spectrum of risky scenarios (including those impossible to mitigate in the first place, see Figure 5), prompting the exploration of training algorithms that can distinguish such cases and exclude them during training. Another interesting future direction from this study is exploring how to design conditional deployment specifications based on the meta-guide model. This will be vital to ensure the safe deployment of robots in evolving real-world scenarios.
@@ VI Conclusion
This research investigates a framework that integrates IS approaches in both the training and evaluation of autonomous agents under complex interactive scenarios. By integrating guided meta reinforcement learning with CE through the IS framework, the training approach adapts efficiently to both common and rare driving conditions. The experimental results using both synthetic and real-world data highlight the potential for substantial improvements in agent performance, particularly in handling interactive T-intersections, with success rates improving significantly across iterative training sessions. Our findings highlight compelling evidence of the efficacy of the integrated IS framework in enhancing the performance of the ego agents over time. Future work will focus on characterizing the diminishing training improvement over time and identifying useful scenarios for training (e.g. by identifying and omitting unavoidable failure cases) to further advance the development of safe autonomous agents.

### v2

@@ IV-A Synthetic Naturalistic Distribution Example
For our initial study, we create an intentional mismatch between naturalistic and training distributions. We set $p_{\text{naturalistic}}=N(1.5,0.5^{2})$ and $p_{\text{training}}=N(0.5,0.5^{2})$, deliberately exposing the ego policy to more challenging scenarios during training than it would encounter in typical operation. In this setup, we implement four training schemes: GEP and GIS trained using uniform distributions, NEP trained on the naturalistic distribution, and CEIS trained using our proposed training distribution ($K=1$, non-iterative). Each policy undergoes extensive testing with 5,000 simulation runs. During evaluation, we sample $\beta$ from a CE-optimized distribution ($p_{\text{evaluation}}$) constructed from isolated test data. We compute final metrics using importance sampling with a naturalistic distribution fitted to the test set and summarize the means and standard deviations of the success rate (SR), collision rate (CR), and timeout rate (TR) in Table I.
We also conduct an ablation study, testing Gaussian training distributions with parameters $\mu\in\{-0.5,0.5,1.5,2.5\}$ and $\sigma\in\{0.5,0.75,1.0\}$, summarized in Table II, to evaluate whether the CE optimized training distribution leads to maximal performance.
TABLE I: Ego policies evaluation for the synthetic dataset
Ego Policy Training Distribution SR $\uparrow$   (avg±$\pm$±std) CR $\downarrow$   (avg±$\pm$±std) TR $\downarrow$   (avg±$\pm$±std)
GEP $U(-1,3)$ 0.66$\pm$0.47 0.33$\pm$0.47 0.01$\pm$0.08
GIS $U(-1,3)$ 0.88$\pm$0.34 0.11$\pm$0.31 0.10$\pm$0.13
NEP $N(1.5,0.5^{2})$ 0.68$\pm$0.47 0.12$\pm$0.14 0.20$\pm$0.32
CEIS $N(0.5,0.5^{2})$ 0.81$\pm$0.39 0.11$\pm$0.39 0.08$\pm$0.31
@@
Std Avg Metric (avg±$\pm$±std) $\mu=-0.5$ $\mu=0.5$ $\mu=1.5$ $\mu=2.5$
$\sigma=0.5$ SR $\uparrow$ 0.76$\pm$0.42 0.81$\pm$0.39 0.68$\pm$0.47 0.70$\pm$0.40
CR $\downarrow$ 0.04$\pm$0.19 0.11$\pm$0.39 0.12$\pm$0.14 0.22$\pm$0.41
TR $\downarrow$ 0.20$\pm$0.11 0.08$\pm$0.31 0.20$\pm$0.32 0.03$\pm$0.08
$\sigma=0.75$ SR $\uparrow$ 0.77$\pm$0.13 0.79$\pm$0.28 0.77$\pm$0.31 0.78$\pm$0.33
CR $\downarrow$ 0.12$\pm$0.15 0.17$\pm$0.25 0.15$\pm$0.46 0.20$\pm$0.31
TR $\downarrow$ 0.10$\pm$0.14 0.05$\pm$0.13 0.08$\pm$0.18 0.02$\pm$0.31
$\sigma=1.0$ SR $\uparrow$ 0.68$\pm$0.49 0.70$\pm$0.39 0.68$\pm$0.47 0.69$\pm$0.31
CR $\downarrow$ 0.22$\pm$0.38 0.19$\pm$0.39 0.21$\pm$0.46 0.19$\pm$0.29
TR $\downarrow$ 0.10$\pm$0.30 0.11$\pm$0.28 0.11$\pm$0.27 0.12$\pm$0.23
@@ IV-B Real-World Naturalistic Distribution
In our second experiment, we evaluate our approach against real-world driving behaviors recorded in the InD dataset [inDdataset] and the RoundD dataset [rounDdataset]. We focus on two T-intersections in Aachen: Heckstrasse and Neukollner Strasse. To extract meaningful behavior parameters, we employ maximum entropy inverse reinforcement learning [huang2021driving] to infer the underlying $\beta$ values from vehicle trajectories. This process assumes our reward structure from Equation (2) and estimates the $\beta$ that best explains each vehicle’s observed behavior, with trajectory likelihood proportional to the exponential of the reward.
@@
Using these naturalistic distributions, we benchmark the same four approaches (GEP, GIS, NEP, and CEIS) as in the synthetic example, with the GEP and GIS using $U(-1,3)$ and NEP with the kernel density estimate from T-Intersection #1 as $p_{\text{training}}$. We run CEIS for $K=10$ iterations and let it discover an interesting progression of behavior distributions, starting with $N(-0.12,0.5^{2})$, then incorporating $N(2.14,0.5^{2})$ and $N(0.81,0.5^{2})$ in the subsequent two iterations.
@@
Figure 3: Naturalistic and CEIS distributions for the RoundD roundabout experiments [rounDdataset].
We extend our analysis to include roundabout scenarios from the rounD dataset [rounDdataset], specifically Neuweilear, Aachen (Roundabout #1) and Thiergarten, Alsdorf (Roundabout #2). The training distributions associated with the roundabout dataset are shown in Figure 3, with CEIS using a Gaussian $N(2.55,0.5^{2})$ for iteration 1 and adding $N(-0.24,0.5^{2})$ and $N(0.97,0.5^{2})$ for iterations 2 and 3, respectively. The complete results of our analysis across all scenarios are presented in Table III and visualized in Figure 4.
TABLE III: Ego policies evaluation for the naturalistic datasets ( InD T-intersections and RoundD roundabouts)
Intersection (avg±$\pm$±std) Roundabout (avg±$\pm$±std)
Ego Policy Training Distribution SR $\uparrow$ CR $\downarrow$ TR $\downarrow$ SR $\uparrow$ CR $\downarrow$ TR $\downarrow$
GEP $U(-1,3)$ 0.75$\pm$0.44 0.23$\pm$0.42 0.01$\pm$0.11 0.79$\pm$0.31 0.15$\pm$0.16 0.05$\pm$0.11
GIS $U(-1,3)$ 0.85$\pm$0.38 0.13$\pm$0.34 0.02$\pm$0.16 0.75$\pm$0.21 0.18$\pm$0.10 0.07$\pm$0.05
NEP KDE 0.82$\pm$0.39 0.08$\pm$0.29 0.10$\pm$0.33 0.80$\pm$0.22 0.03$\pm$0.06 0.16$\pm$0.09
CEIS, $k=1$ $N(\mu,\sigma^{2})$ 0.83$\pm$0.38 0.11$\pm$0.32 0.06$\pm$0.26 0.73$\pm$0.15 0.15$\pm$0.06 0.12$\pm$0.05
CEIS, $k=2$ GMM 0.89$\pm$0.32 0.08$\pm$0.29 0.03$\pm$0.20 0.78$\pm$0.18 0.12$\pm$0.07 0.10$\pm$0.04
CEIS, $k=3$ GMM 0.93$\pm$0.28 0.03$\pm$0.19 0.03$\pm$0.21 0.82$\pm$0.16 0.08$\pm$0.04 0.08$\pm$0.05
@@ V-A Dominant IS-based Performance Over the Baselines
In both synthetic and real-world experiments, IS-based algorithms demonstrate statistically significant performance advantages over their non-IS counterparts (see Table I and Table III). In synthetic scenarios, GIS achieves 88% success rate and 11% collision rate, while CEIS achieves 81.1% success rate and 10.8% collision rate—representing over 10% improvement compared to non-IS approaches. The baseline methods show notably weaker performance, with NEP achieving only 68 and GEP 66% success rates, along with concerning collision rates of 12% and 33%, respectively.
This performance gap becomes even more pronounced in real-world scenarios, where CEIS achieves a remarkable 93% success rate in intersections and 82% in roundabouts by iteration 3 (and 98% and 94% by iteration 10, respectively), while maintaining minimal collision rates of 3% (see Figure 4). The only trade-off appears in timeout rates, where GEP shows marginally better performance with 0.6% timeouts. However, we argue that in practical robotics applications, ensuring safety through careful risk assessment outweighs the marginal benefits of faster task completion. The IS-based approach demonstrates superior risk reasoning by explicitly considering the likelihood of dangerous maneuvers in its decision-making process (as visualized in Figure 5). This represents a key strength of our framework: by systematically evaluating performance across both common and rare scenarios, IS enables more robust adaptation to real-world driving dynamics. The result is an agent that not only excels in typical situations but also maintains reliable performance in unexpected or challenging conditions, as evidenced by the consistently lower standard deviations in CEIS (34%) compared to GEP (44%).
@@ V-B Better Results in Naturalistic Settings
While GIS shows stronger performance in synthetic scenarios (88.0% success rate), CEIS demonstrates superior adaptability in naturalistic environments. This is particularly evident in its consistent performance improvements across different scenarios types, achieving a 93% success rate in intersections and 82% in roundabouts by iteration 3, while GIS slightly underperforms at 85% and 75%, respectively. The sensitivity analysis in Table II reinforces this finding, showing optimal performance with concentrated training distributions ($\sigma=0.5$), especially when combined with CEIS’s iterative refinement strategy.
This observation suggests that CEIS’s iterative training scheme enables effective continuous learning and adaptation. However, it is important to note that our current implementation focuses on single-parameter behavioral models ($\beta$), which may not fully capture the complexity of multi-agent interactions in higher-dimensional spaces. However, scaling to more complex scenarios involving multiple types of road users or varying weather conditions would require more sophisticated behavioral models than the current T-intersection and roundabout models used in this study, and thus may require additional research. Nevertheless, these current results demonstrate CEIS’ potential as a framework for developing safe and more balanced ego policy to deal with highly interactive environments.
((a)) Success (the rear vehicle given sufficient reaction time)
((b)) Failure (merging initiated too quickly causing a crash)
((c)) Success (the ego safely navigating the roundabout)
((d)) Failure (overly slow merging causing a rear end)
Figure 5: Success and failure examples of the ego vehicle policy in the T-intersection and roundabout scenarios.
V-C Trend of Improving Performance Over Time
A longitudinal evaluation of CEIS (Figure 4) reveals interesting scenario-dependent learning dynamics. In T-intersections, CEIS demonstrates rapid initial improvement—success rates climb sharply from 83% to 93% within just three iterations, with an impressive collision rate reduction from 11% to 3%. The learning curve for roundabouts, while positive, shows more modest gains from 73% to 82% over the same period, with collision rates decreasing from 15% to 8%. This different learning rates suggest important insights about scenario complexity. T-intersections show approximately 5% improvement per iteration in the first three iterations, compared to 3.4% for roundabouts. Moreover, the timeout rates exhibit particularly interesting behavior: they converge much faster in T-intersections (dropping below 3% by iteration 2) than in roundabouts (requiring 4-5 or even more iterations), suggesting different risk-learning dynamics and that CEIS learns to balance safety and progress more efficiently in simpler geometric configurations.
More importantly, however, a critical limitation emerges: both scenarios eventually hit performance plateaus (98% vs 94% projected success rates), albeit at different iteration counts. This suggests that CEIS’s current formulation might struggle with scenario-specific ‘irreducible complexity’,i.e. cases where the inherent multi-agent interaction patterns create unavoidable failure modes (see Figure 6 for two such examples from the intersection experiments). This observation opens exciting research directions, particularly in developing scenario-aware training strategies that could dynamically adjust sampling distributions based on detected complexity levels, potentially breaking through these scenario-specific performance ceilings.
V-D Generalization and Scalability
Our framework demonstrates robust generalization capabilities across diverse driving scenarios, supported by concrete performance metrics. In synthetic environments, CEIS achieves 81% success rate, while in real-world settings it reaches 98% for intersections and 94% for roundabouts. This consistent performance spans not only across scenario types but also across behavioral distributions—from simplified synthetic models to complex real-world patterns extracted from the InD and RounD datasets, suggesting effective transfer of learned safety behaviors across domains.
However, scalability presents notable challenges. While CEIS shows computational efficiency in current scenarios, scaling to higher-dimensional problems faces two key limitations: the growing computational cost of CE optimization and the increasing complexity of GMM fitting in larger parameter spaces. Additionally, extending to multi-modal sensing scenarios (cameras, LiDARs, IMUs) would require sophisticated noise modeling and distribution estimation techniques. These scalability challenges point to promising research directions, particularly in developing efficient alternatives to CE optimization [arief2021deep, deo2023achieving] and exploring adaptive GMM architectures for high-dimensional spaces. Future work might also investigate transfer learning approaches to reduce the computational burden when adapting to new sensor modalities or scenario types.
((a)) Sudden acceleration.
((b)) Overcrowded scene.
@@ VI Conclusion
This research investigates a framework that integrates IS approaches in both the training and evaluation of autonomous agents under complex interactive scenarios. By integrating guided meta reinforcement learning with CE through the IS framework, the training approach adapts efficiently to both common and rare driving conditions. The experimental results using both synthetic and real-world data highlight the potential for substantial improvements in agent performance, with success rates improving significantly across iterative training sessions. Our findings highlight compelling evidence of the potential to integrate IS framework in both training and evaluation to enhance the performance of the ego agents over time. Future work will focus on exploring the scalability of the proposed framework to more complex multi-agent scenarios and sensor modalities, as well as developing adaptive training strategies to address scenario-specific performance ceilings.